OpenStudy (anonymous):

MEDAL AND FAN!!!!<3 a rectangle has an area 264cm. and length of 22cm. what is the width of a rectangle. a.12cm b.22cm. c.24cm d.44cm

OpenStudy (anonymous):

A=w*l

OpenStudy (smileyxl3):

W=A/L

OpenStudy (the_fizicx99):

A = L * W Since you don't have your W you divide, W = A/L W = 264/22

OpenStudy (anonymous):

so 12cm is my answer

OpenStudy (the_fizicx99):

Yes

OpenStudy (the_fizicx99):

12*22 = 264